This is a word problem involving ratios that I have been struggling with for a couple of hours now. Can you please help me.
The ratios of teachers to students at Elaine's school is 4 to 52. If there are 27 teachers at this school, then how many students are there? write a complete solution.
This what I have been trying: 4/52 = x/27, cross multiplied and 52x = 108, but it is not coming out right.

Answer by Earlsdon(6294)

You have the right idea but the second ratio of your proportion is inverted.
Remember to keep like-items in the tops and like-items in the bottoms of the ratios.
Teachers/students = 4/52 You had this correctly shown.
Let x = # of students, so you want to see the number of teachers on top and the number of students (x) on the bottom.
27/x
The proportion looks like:
4/52 = 27/x Solve for x by cross-multiplying.
4x = (52)(27) Divide both sides by 4.
x = 1404/4
x = 351 This is the number of students.
Check:
4/52 = 27/351 Reduce the fractions.
1/13 = 1/13